Required information

Problem 14-15 Comprehensive Ratio Analysis [LO14-2, LO14-3, LO14-4, LO14-5, LO14-6]

[The following information applies to the questions displayed below.]

You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked you to perform a comprehensive analysis of the company’s financial statements, including comparing Lydex’s performance to its major competitors. The company’s financial statements for the last two years are as follows:

Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 870,000 $ 1,110,000
Marketable securities 0 300,000
Accounts receivable, net 2,340,000 1,440,000
Inventory 3,510,000 2,100,000
Prepaid expenses 240,000 180,000
Total current assets 6,960,000 5,130,000
Plant and equipment, net 9,340,000 8,960,000
Total assets $ 16,300,000 $ 14,090,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 3,920,000 $ 2,800,000
Note payable, 10% 3,600,000 3,000,000
Total liabilities 7,520,000 5,800,000
Stockholders' equity:
Common stock, $70 par value 7,000,000 7,000,000
Retained earnings 1,780,000 1,290,000
Total stockholders' equity 8,780,000 8,290,000
Total liabilities and stockholders' equity $ 16,300,000 $ 14,090,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,770,000 $ 12,680,000
Cost of goods sold 12,616,000 9,510,000
Gross margin 3,154,000 3,170,000
Selling and administrative expenses 1,594,000 1,568,000
Net operating income 1,560,000 1,602,000
Interest expense 360,000 300,000
Net income before taxes 1,200,000 1,302,000
Income taxes (30%) 360,000 390,600
Net income 840,000 911,400
Common dividends 350,000 455,700
Net income retained 490,000 455,700
Beginning retained earnings 1,290,000 834,300
Ending retained earnings $ 1,780,000 $ 1,290,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.3
Acid-test ratio 1.0
Average collection period 30 days
Average sale period 60 days
Return on assets 8.3 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.9
Price-earnings ratio 10

rev: 04_27_2020_QC_CS-209476

Problem 14-15 Part 3

3. You decide, finally, to assess the company’s liquidity and asset management. For both this year and last year, compute:

a. Working capital.

b. The current ratio. (Round your final answers to 2 decimal places.)

c. The acid-test ratio. (Round your final answers to 2 decimal places.)

d. The average collection period. (The accounts receivable at the beginning of last year totaled $1,580,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

e. The average sale period. (The inventory at the beginning of last year totaled $1,940,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

f. The operating cycle. (Round your intermediate calculations and final answers to 2 decimal place.)

g. The total asset turnover. (The total assets at the beginning of last year totaled $12,980,000.) (Round your final answers to 2 decimal places.)
Answer #1

a. Working Capital

Working Capital = Current Assets - Current Liabilities
Particulars This Year Last Year
Current Assets $6,960,000 $5,130,000
Less: Current Liabilities $3,920,000 $2,800,000
Working Capital $3,040,000 $2,330,000

b. Current Ratio

Current Ratio = Current Assets / Current Liabilities
Particulars This Year Last Year
Current Assets $6,960,000 $5,130,000
/ Current Liabilities $3,920,000 $2,800,000
Current Ratio 1.78 : 1 1.83 : 1

c. Acid-test ratio

Acid Test Ratio = Quick Assets / Current Liabilities

Quick Assets = Current Assets - Inventory - Prepaid expenses

Particulars This Year Last Year
Current Assets $6,960,000 $5,130,000
Less: Inventory $3,510,000 $2,100,000
Less: Prepaid expenses $240,000 $180,000
Quick Assets $3,210,000 $2,850,000
/ Current Liabilities $3,920,000 $2,800,000
Acid Test Ratio 0.82 : 1 1.02 : 1

d. Average collection period

Average collection period = 365 days / Credit Sales * Average accounts receivable

This Year

Average accounts receivable = ($1,440,000 + $2,340,000) / 2 = $1,890,000

Average collection period = 365 days / $15,770,000 * $1,890,000 = 43.74 days

Last Year

Average accounts receivable = ($1,580,000 + $1,440,000) / 2 = $1,510,000

Average collection period = 365 days / $12,680,000 * $1,510,000 = 43.47 days

e. Average sale period

Average sale period = 365 days / Cost of goods sold * Average Inventories

This Year

Average Inventories = ($2,100,000 + $3,510,000) / 2 = $2,805,000

Average sale period = 365 days / $12,616,000 * $2,805,000 = 81.15 days

Last Year

Average Inventories = ($1,940,000 + $2,100,000) / 2 = $2,020,000

Average sale period = 365 days / $9,510,000 * $2,020,000 = 77.53 days

f. Operating Cycle

Operating cycle = Average collection period + Average sale period

This Year

Operating Cycle = 43.74 days + 81.15 days = 124.89 days

Last Year

Operating Cycle = 43.47 days + 77.53 days = 121.00 days

g. Total asset turnover

Asset turnover = Credit Sales / Average total assets

This Year

Average total assets = ($14,090,000 + $16,300,000) / 2 = $15,195,000

Asset Turnover = $15,770,000 / $15,195,000 = 1.04 times

Last Year

Average total assets = ($12,980,000 + $14,090,000) / 2 = $13,535,000

Asset Turnover = $12,680,000 / $13,535,000 = 0.94 times

All the best...